OpenStudy (anonymous):

the area of a circle of radius r is the same as the area of a square of side a.Ratio of r to a is?

OpenStudy (anonymous):

do u know, what formula to find the area of a cicle and a square

OpenStudy (anonymous):

area of a circle = pi x r x r = pi*r^2 area of a square = a x a = a^2

OpenStudy (anonymous):

but what will be the ratio @tanjung